# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v2

@@ A.1. Comparison Principle
Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63). If $F,G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and if $M$ is the constant defined by
$M=\lim_{\delta\downarrow 0}\sup\left\{(F(x,0)-G(y,0))_{+}\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\},$
@@
(66) $\sup\left\{F(x,t)-G(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq Me^{Lt}$
In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (61) for some fixed initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
@@
The comparison principle implies, in particular, that there is at most one viscosity solution of (61) if $F_{\text{in}}$ is continuous. It is known that, indeed, a solution exists in this case, and the assumption (64) ensures that, for each $t>0$, $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$.
@@
Assume that $\mathcal{F}$ and $Q$ satisfy (62), (63), and (64). For any continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, there is a unique viscosity solution $F$ of (61). Further, for any $t>0$, the function $F(\cdot,t)$ is in $CDF(\overline{\mathbb{R}})$.
@@
Existence of a solution taking values in $[0,1]$ follows from Perron’s Method, which is applicable since the constant functions 0 and $1$ are, respectively, viscosity sub- and supersolutions by (64). By the comparison principle, the solution is unique and it takes values in $[0,1]$. Finally, for any $y\in\mathbb{R}$, the function $(x,t)\mapsto F(x+y,t)$ is a viscosity solution of (61) with the initial datum $x\mapsto F_{\text{in}}(x+y)$. Thus, since $F_{\text{in}}$ is nondecreasing, by the comparison principle, $F(x+y,t)\leq F(x,t)$ for all $t\geq 0$ if $y\geq 0$. Since $F$ is continuous, this proves $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$ for each $t>0$. ∎
@@
Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63). Let $\underline{F}$ and $\overline{F}$ be the maximal subsolution and minimal supersolution of (61), respectively, for some given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,\infty)$, then any discontinuous viscosity solution $F$ of (61) must also satisfy $F=\underline{F}$ in $\mathbb{R}\times(0,\infty)$. In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (61), which is continuous in $\mathbb{R}\times(0,+\infty)$.
@@
As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\underline{F}\leq F\leq\overline{F}$ in $\mathbb{R}\times[0,\infty)$ in general. Therefore, in the present setting, $F=\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, implying that there is at most one discontinuous viscosity solution (up to arbitrary choices at $t=0$ at points where $F_{\text{in}}$ is discontinuous).
The fact that at least one exists is almost immediate: Since $\underline{F}$ is a supremum of subsolutions, the upper semicontinuous envelope $\underline{F}^{*}$ satisfies $\partial_{t}\underline{F}^{*}-\mathcal{F}(\partial_{x}\underline{F}^{*},\partial_{x}^{2}\underline{F}^{*})+Q(\underline{F}^{*})\leq 0$ in $\mathbb{R}\times(0,+\infty)$, and solutions of (61) with initial data larger than $F_{\text{in}}$ can be used to prove $\underline{F}^{*}(\cdot,0)=F_{\text{in}}$ at $t=0$. Therefore, by comparison, $\underline{F}^{*}\leq\overline{F}$ in $\mathbb{R}\times[0,+\infty)$. A similar argument shows that $\partial_{t}\overline{F}_{*}-\mathcal{F}(\partial_{x}\overline{F}_{*},\partial^{2}_{x}\overline{F}_{*})+Q(\overline{F}_{*})\geq 0$ in $\mathbb{R}\times(0,+\infty)$ and $\overline{F}_{*}(\cdot,0)$ equals the lower semicontinuous envelope of $F_{\text{in}}$, so $\overline{F}_{*}\geq\underline{F}$. Since $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$ and the upper (resp. lower) semicontinuous envelope of a function is always larger (resp. smaller) than the function itself, this proves $\underline{F}$ is continuous in $\mathbb{R}\times(0,+\infty)$ and a discontinuous viscosity solution of (61). ∎
@@
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. Suppose that $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ are two sequences of continuous functions in $CDF(\overline{\mathbb{R}})$ such that
@@ A.2. Uniqueness for Nondecreasing Data
In this section, the PDE’s of interest here are shown to satisfy the uniqueness criterion of Proposition 13. For the Hamilton-Jacobi equations arising in the $\text{Beta}(2,1)$ case, the fact that almost-everywhere-continuous data leads to a unique solution at positive times is known from [11]. In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows, for instance, from the characterization of nonnegative solutions of the porous medium equation. Short, simplified proofs are presented here for the reader’s convenience.
@@
(69) $\lim_{|p|\to\infty}|p|^{-1}H(p)=+\infty.$
At this stage, it is worth noting that there is no loss of generality in assuming that $H$ is convex. Indeed, in general, if $\partial_{t}F+H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$, then the function G(x,t)=1F(x,t)G(x,t)=1-F(-x,t) satisfies
$\partial_{t}G-H(-\partial_{x}G)=0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,\infty)$
and $H$ is convex if and only if the function $p\mapsto-H(-p)$ is concave.
For superlinear convex/concave $H$, the Hopf-Lax formula furnishes the unique discontinuous viscosity solution of (61).
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$. If $\mathcal{F}(v,w)=-H(v)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying $H(0)=0$ and (69), and if $\underline{F}$ and $\overline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (61), then, for any $(x,t)\in\mathbb{R}\times(0,\infty)$,
@@
Some regularity of $F_{\text{in}}$ is necessary to ensure that $\overline{F}$ and $\underline{F}$ coincide at positive times. For instance, $F_{\text{in}}=\mathbf{1}_{\mathbb{Q}}$ provides a counterexample. An optimal version of the result can be found in [11].
@@
Therefore, since $F_{\text{in}}$ is right-continuous, this proves $\lim_{N\to\infty}F^{(N),+}\leq F$ pointwise in $\mathbb{R}\times(0,+\infty)$. Next, by superlinearity of $L$, for any $(x,t)$, it is possible to write $F^{(N),-}(x,t)=F^{(N),-}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$. If $y_{N}\to y$ as $N\to\infty$ (say, along a subsequence), then
@@
Notice that if $F$ were a smooth solution of $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=0$, then the derivative $\rho=\partial_{x}F$ would solve the porous medium equation (PME) $\partial_{t}\rho=\frac{a}{m}\partial_{x}^{2}(\rho^{m})$. To prove the uniqueness of solutions with possibly discontinuous, nondecreasing data, this reasoning will be inverted.
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$. If $\mathcal{F}$ has the form (71) for some $m>1$, then the maximal subsolution $\underline{F}$ and minimal supersolution $\overline{F}$ of (61) are given for $(x,t)\in\mathbb{R}\times(0,\infty)$ by the formula
@@
(75) $\left\{\begin{array}[]{r l}\partial_{t}\rho-\frac{a}{m}\partial_{x}^{2}(\rho^{m})=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ \rho(0)=\partial_{x}F_{\text{in}}.\end{array}\right$.
@@
The uniqueness and claimed representation formula for the solution follows from the classical uniqueness theorem for the PME by Dahlberg and Kenig [16]. To see this, first, note that if $\partial_{x}F_{\text{in}}$ is smooth and positive, then it is well-known that the solution $\rho$ of the PME is smooth (up to time $t=0$), hence the antiderivative given by (72) is a smooth solution of (61). It is classical that smooth solutions are viscosity solutions. It is also classical in this setting that $\int_{-\infty}^{\infty}\rho(y,t)\,dy=F_{\text{in}}(+\infty)-F_{\text{in}}(-\infty)$ for each $t>0$, as in item (i).
@@
Finally, if $\partial_{x}F_{\text{in}}$ is merely a finite measure, consider sequences of smooth functions $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ as in Lemma 1, chosen in such a way that $F^{(N),\pm}_{\text{in}}(\bar{x})\to F_{\text{in}}(\bar{x})$ as $N\to+\infty$ for each $\bar{x}\in\{-\infty,+\infty\}$. By the bounds (76), the corresponding solutions of the PME $\{\rho^{(N),+}\}$ and $\{\rho^{(N),-}\}$ started from the data $\{\partial_{x}F^{(N),+}_{\text{in}}\}$ and $\{\partial_{x}F^{(N),-}\}$ are uniformly bounded and equicontinuous in $\mathbb{R}\times(\tau,+\infty)$ for any $\tau>0$. Therefore, the antiderivatives $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ become equicontinuous after a positive amount of time. Since they are, respectively, monotone decreasing and monotone increasing in $N$, the limits $F^{+}=\lim_{N\to\infty}F^{(N),+}$ and $F^{-}=\lim_{N\to\infty}F^{(N),-}$ exist. Since $F^{(N),-}\leq F^{-}\leq F^{+}\leq F^{(N),+}$ holds in $\mathbb{R}\times[0,+\infty)$ for any $N$, $F^{-}(\cdot,t)$ and $F^{+}(\cdot,t)$ both converge vaguely to $F_{\text{in}}$ as $t\downarrow 0$. By Arzelà-Ascoli, the derivatives $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ are continuous distributional solutions of the PME in $\mathbb{R}\times(0,+\infty)$, and the vague convergence of $F^{-}$ and $F^{+}$ at time $t=0$ implies that the initial trace of $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ equals $\partial_{x}F_{\text{in}}$. Therefore, by Dahlberg-Kenig, $\partial_{x}F^{-}=\partial_{x}F^{+}$ in $\mathbb{R}\times(0,+\infty)$, which implies the difference $F^{+}-F^{-}$ depends only on time.
@@
The identity $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,\infty)$ still holds for the initial-value problem (61) with arbitrary data in $CDF(\overline{\mathbb{R}})$ if $\mathcal{F}$ satisfies the assumptions of Propositions 14 or 15 and $Q$ satisfies (63) and (64).
@@
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $F_{\text{in}}$ is continuous, then Proposition 12 implies there is a unique viscosity solution $F$ of (61) and $\overline{F}=\underline{F}=F$. Otherwise, if it has jumps, fix sequences $\{F_{\text{in}}^{(N),+}\}$ and $\{F_{\text{in}}^{(N),-}\}$ of continuous functions in $CDF(\overline{\mathbb{R}})$ satisfying (67) and (68) as in Lemma 1 and, in addition,
@@
$\limsup_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(B)\delta e^{L(t-\delta)}.$

### v3

@@ A.1. Comparison Principle
Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63). If $F,G:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and if $M$ is the constant defined by
$M=\lim_{\delta\downarrow 0}\sup\left\{(F(x,0)-G(y,0))_{+}\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|x-y|\leq\delta\right\},$
@@
(66) $\sup\left\{F(x,t)-G(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq Me^{Lt}.$
In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (61) for some fixed initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,+\infty)$.
@@
The comparison principle implies, in particular, that there is at most one viscosity solution of (61) if $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ is continuous. It is known that, indeed, a solution exists in this case, and the assumption (64) ensures that, for each $t>0$, $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$.
@@
Assume that $\mathcal{F}$ and $Q$ satisfy (62), (63), and (64). For any continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, there is a unique bounded (continuous) viscosity solution $F$ of (61). Further, for any $t>0$, the function $F(\cdot,t)$ is in $CDF(\overline{\mathbb{R}})$.
@@
The existence of a bounded solution follows from Perron’s Method, see [15] and the references therein. By Theorem 4, this is the unique bounded solution. By (64), since $0\leq F_{\text{in}}\leq 1$, the constant functions $G(x,t)\equiv 1$ and $G(x,t)\equiv 0$ are respectively viscosity super- and subsolutions of (61), which implies that $0\leq F(x,t)\leq 1$ for any $(x,t)$ by comparison. Finally, for any $y\in\mathbb{R}$, the function $(x,t)\mapsto F(x+y,t)$ is a viscosity solution of (61) with the initial datum $x\mapsto F_{\text{in}}(x+y)$. Thus, since $F_{\text{in}}$ is nondecreasing, the comparison principle implies that $F(x+y,t)\leq F(x,t)$ for all $t\geq 0$ if $y\geq 0$. Altogether, this proves that $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$ for each $t>0$. ∎
@@
Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63). Let $\underline{F}$ and $\overline{F}$ be the maximal subsolution and minimal supersolution of (61), respectively, for some given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, then any discontinuous viscosity solution $F$ of (61) must also satisfy $F=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$. In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (61), which is continuous in $\mathbb{R}\times(0,+\infty)$.
@@
As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\underline{F}\leq F\leq\overline{F}$ in $\mathbb{R}\times[0,+\infty)$ in general. Therefore, in the present setting, $F=\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, implying that there is at most one discontinuous viscosity solution (up to arbitrary choices at $t=0$ at points where $F_{\text{in}}$ jumps).
The fact that at least one exists is almost immediate: Since $\underline{F}$ is a supremum of subsolutions, the upper semicontinuous envelope $\underline{F}^{*}$ satisfies $\partial_{t}\underline{F}^{*}-\mathcal{F}(\partial_{x}\underline{F}^{*},\partial_{x}^{2}\underline{F}^{*})+Q(\underline{F}^{*})\leq 0$ in $\mathbb{R}\times(0,+\infty)$, and solutions of (61) with initial data larger than $F_{\text{in}}$ can be used to prove $\underline{F}^{*}(\cdot,0)=F_{\text{in}}$ at $t=0$. Therefore, by comparison, $\underline{F}^{*}\leq\overline{F}$ in $\mathbb{R}\times[0,+\infty)$. A similar argument shows that $\partial_{t}\overline{F}_{*}-\mathcal{F}(\partial_{x}\overline{F}_{*},\partial^{2}_{x}\overline{F}_{*})+Q(\overline{F}_{*})\geq 0$ in $\mathbb{R}\times(0,+\infty)$ and $\overline{F}_{*}(\cdot,0)$ equals the lower semicontinuous envelope of $F_{\text{in}}$, so $\overline{F}_{*}\geq\underline{F}$. Since $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$ by assumption and the upper (resp. lower) semicontinuous envelope of a function is always larger (resp. smaller) than the function itself, this proves $\underline{F}$ is continuous in $\mathbb{R}\times(0,+\infty)$ and a discontinuous viscosity solution of (61). ∎
@@
Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63) and fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. Suppose that $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ are two sequences of continuous functions in $CDF(\overline{\mathbb{R}})$ such that
@@ A.2. Uniqueness for Nondecreasing Data
In this section, the PDE’s of interest here are shown to satisfy the uniqueness criterion of Proposition 13. For the Hamilton-Jacobi equations arising in the $\text{Beta}(2,1)$ case, the fact that almost-everywhere-continuous data leads to a unique solution at positive times is known from [11]. In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows, for instance, from the characterization of nonnegative solutions of the porous medium equation. Short proofs are presented here for the reader’s convenience.
@@
(69) $\lim_{|v|\to\infty}|v|^{-1}|H(v)|=+\infty.$
At this stage, it is worth noting that there is no loss of generality in assuming that $H$ is convex. Indeed, in general, if $\partial_{t}F+H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$, then the function G(x,t)=1F(x,t)G(x,t)=1-F(-x,t) satisfies
$\partial_{t}G-H(\partial_{x}G)=0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,+\infty)$
and $H$ is convex if and only if the function $v\mapsto-H(v)$ is concave.
For superlinear convex/concave $H$, the Hopf-Lax formula furnishes the unique discontinuous viscosity solution of (61) when $Q=0$.
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$. If $\mathcal{F}(v,w)=-H(v)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying $H(0)=0$ and (69), and if $\underline{F}$ and $\overline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (61), then, for any $(x,t)\in\mathbb{R}\times(0,+\infty)$,
@@
Some regularity of $F_{\text{in}}$ is necessary to ensure that $\underline{F}$ and $\overline{F}$ coincide at positive times. For instance, $F_{\text{in}}=\mathbf{1}_{\mathbb{Q}}$ provides a counterexample. An optimal version of the result can be found in [11].
@@
Therefore, since $F_{\text{in}}$ is right-continuous, this proves $\lim_{N\to\infty}F^{(N),+}\leq F$ pointwise in $\mathbb{R}\times(0,+\infty)$. Next, since the superlinearity of $H$ implies that $L$ is superlinear (see [18, Theorem 3, Section 3.3.2]), for any $(x,t)$, it is possible to write $F^{(N),-}(x,t)=F^{(N),-}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$. If $y_{N}\to y$ as $N\to\infty$ (say, along a subsequence), then
@@
Notice that if $F$ were a smooth strictly increasing solution of $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=0$, then the derivative $\rho=\partial_{x}F$ would solve the porous medium equation (PME) $\partial_{t}\rho=\frac{a}{m}\partial_{x}^{2}(\rho^{m})$. To prove the uniqueness of solutions with possibly discontinuous, nondecreasing data, this reasoning will be inverted.
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$. If $\mathcal{F}$ has the form (71) for some $m>1$, then the maximal subsolution $\underline{F}$ and minimal supersolution $\overline{F}$ of (61) are given for $(x,t)\in\mathbb{R}\times(0,+\infty)$ by the formula
@@
(75) $\left\{\begin{array}[]{r l}\partial_{t}\rho-\frac{a}{m}\partial_{x}^{2}(\rho^{m})=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ \rho(0)=\partial_{x}F_{\text{in}}.\end{array}\right$.
@@
The uniqueness and claimed representation formula for the solution follows from the classical uniqueness theorem for the PME by Dahlberg and Kenig [16]. To see this, first, note that if $\partial_{x}F_{\text{in}}$ is smooth and positive, then it is well-known that the solution $\rho$ of the PME is smooth, hence the antiderivative given by (72) is a smooth solution of (61). It is classical that smooth solutions are viscosity solutions. It is also classical in this setting that $\int_{-\infty}^{\infty}\rho(y,t)\,dy=F_{\text{in}}(+\infty)-F_{\text{in}}(-\infty)$ for each $t>0$, as in item (i).
@@
Finally, if $\partial_{x}F_{\text{in}}$ is merely a finite measure, consider sequences of smooth functions $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ as in Lemma 1, chosen in such a way that $F^{(N),\pm}_{\text{in}}(\bar{x})\to F_{\text{in}}(\bar{x})$ as $N\to+\infty$ for each $\bar{x}\in\{-\infty,+\infty\}$. By the bounds (76), the corresponding solutions of the PME $\{\rho^{(N),+}\}$ and $\{\rho^{(N),-}\}$ started from the data $\{\partial_{x}F^{(N),+}_{\text{in}}\}$ and $\{\partial_{x}F^{(N),-}\}$ are uniformly bounded and equicontinuous in $\mathbb{R}\times(\tau,+\infty)$ for any $\tau>0$. Therefore, the antiderivatives $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ become equicontinuous after a positive amount of time. Since they are, respectively, monotone decreasing and monotone increasing in $N$, the limits $F^{+}=\lim_{N\to\infty}F^{(N),+}$ and $F^{-}=\lim_{N\to\infty}F^{(N),-}$ exist. Since $F^{(N),-}\leq F^{-}\leq F^{+}\leq F^{(N),+}$ holds in $\mathbb{R}\times[0,+\infty)$ for any $N$, $F^{-}(\cdot,t)$ and $F^{+}(\cdot,t)$ both converge vaguely to $F_{\text{in}}$ as $t\downarrow 0$. By Arzelà-Ascoli, the derivatives $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ are continuous distributional solutions of the PME in $\mathbb{R}\times(0,+\infty)$, and the vague convergence of $F^{-}$ and $F^{+}$ at time $t=0$ implies that the initial traces of $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ both equal $\partial_{x}F_{\text{in}}$. Therefore, by Dahlberg-Kenig, $\partial_{x}F^{-}=\partial_{x}F^{+}$ in $\mathbb{R}\times(0,+\infty)$, which implies the difference $F^{+}-F^{-}$ depends only on time.
@@
The identity $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$ still holds for the initial-value problem (61) with arbitrary data in $CDF(\overline{\mathbb{R}})$ if $\mathcal{F}$ satisfies the assumptions of Propositions 14 or 15 and $Q$ satisfies (63) and (64).
@@
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $F_{\text{in}}$ is continuous, then Proposition 12 implies there is a unique viscosity solution $F$ of (61) and $\underline{F}=\overline{F}=F$. Otherwise, if it has jumps, fix sequences $\{F_{\text{in}}^{(N),+}\}$ and $\{F_{\text{in}}^{(N),-}\}$ of continuous functions in $CDF(\overline{\mathbb{R}})$ satisfying (67) and (68) as in Lemma 1 and, in addition,
@@
$\limsup_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(Q)\delta e^{L(t-\delta)}.$